Find the equation in point slope form of the line that is perpendicular inspector of the segment between (16,-4) and (-2,-76)

Answer by lwsshak3(11628)

Find the equation in point slope form of the line that is the perpendicular bisector of the segment between (16,-4) and (-2,-76)
**
I am assuming you meant perpendicular bisector for perpendicular inspector:
..
slope of line connecting given points=∆y/∆x=(-76-(-4))/(-2-16)=-72/-18=4
coordinates of perpendicular bisector=midpoint of given points=((x1+x2)/2, (y1+y2)/2)=(16-2)/2,(-4-76)/2=(7,-40)
slope of line perpendicular to given line=-1/4 (negative reciprocal)
form of equation for a line: y=mx+b, m=slope, b=y-intercept
y=-x/3+b
solve for b using coordinates of perpendicular bisector
-40=-7/4+b
b=-153/4
equation: y=-x/4-153/4